# Wasserstein Distributionally Robust Shallow Convex Neural Networks

### v2

@@ 8.2.2 Numerical Results
We observe that the WaDiRo-SCNN performs on par with the GP as both models seem adequate to model nonlinear patterns. Still, its performance is slightly above as it has the lower MAE between the two on 10 buildings out of 16 while having similar RMSEs. Note that trained GPs may only converge to local optima. We observe that the linear SVR has the worst performance of all models as it has a consistently lower performance than the WaDiRo linear regression. We observe no significant performance gap between the unconstrained and physics-constrained WaDiRo-SCNNs. We also remark, as shown in Table 3, that the WaDiRo-PCSCNN has the lowest number of constraint violations throughout the experiment and that the WaDiRo-SCNN has the second lowest amount with yet 59.12% more. Additional figures, viz., the absolute errors and test prediction plots, are presented in Appendix C for completeness.
@@ 8.3.2 Numerical results
Experiment results are presented in Figures 5 and 6. We observe some general tendencies across the two datasets and the different models. The minimum and maximum of each scale are identified in the top left corner of each subfigure. First, the lowest test errors are achieved for high neuron counts and low regularization. Instability seems inversely correlated with the size of the regularization parameter. Thus, there seems to be a tradeoff between precision and stability. We note that some neurons count are anomalously less stable than others, e.g., 6 neurons in Figure 5, we believe that this is related to the population of sampling vectors used as the same outlying patterns emerge similarly for all models.
The experiment on the energy dataset demonstrates that neural networks can be highly unstable with lower regularization. Indeed, with the added effect of the scaling, $L_{\varepsilon}$ reaches the millions. For a high number of neurons and an intermediate regularization parameter, we note that there seems to be a sweet spot where test errors are low and stability is high.
@@ 9 Closing remarks
Acknowledgments
This work was possible thanks to funding from the Fonds de recherche du Québec (FRQ), the Natural Sciences and Engineering Research Council of Canada (NSERC), Mitacs Accelerate, and Hilo by Hydro-Québec. Special thanks to Salma Naccache and Bertrand Scherrer for their active support and the enriching discussions, as well as Steve Boursiquot, Ahmed Abdellatif, and Odile Noël from Hilo for making this project possible.
@@ A Proof of Proposition 1
$h_{\bm{\beta}}^{*}(\bm{\theta})<+\infty\iff\sup_{\hat{\mathbf{z}}:\hat{\mathbf{z}}^{\top}\bm{\beta}\geq 0}\{\bm{\theta}^{\top}\hat{\mathbf{z}}-{\bm{\beta}}^{\top}\hat{\mathbf{z}}\}<+\infty\ \ \text{and}\ \sup_{\hat{\mathbf{z}}:\hat{\mathbf{z}}^{\top}\bm{\beta}\leq 0}\{\bm{\theta}^{\top}\hat{\mathbf{z}}+{\bm{\beta}}^{\top}\hat{\mathbf{z}}\}<+\infty.$
@@
[…]
@@
[…]
@@
[…]
@@
$\min_{\bm{\nu},\bm{\omega}}\quad$ $\epsilon a$ $+\frac{1}{N}\sum_{j=1}^{N}c_{j}$
@@
$-{\bm{\beta}}_{i}$ $\leq a\quad\forall i\in\llbracket 2Pd+1\rrbracket$
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\bm{\beta}}}$ ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}=\left(\operatorname{vec}(\bm{\nu}),\operatorname{vec}(-\bm{\omega}),-1\right)}}$
@@ B.1 Hyperparameters and setting
{tblr}
colspec=m3cm—lll,hline1,8 = 1.5pt, hline2,3,4,5,6,7=1pt Model & Hyperparameter Description Possible values
@@
Norm of the Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Regularization framework $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Norm of the Wasserstein metric $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Regularization framework $[\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
dropout_p
Size of each batch
@@
We sample 2000 data points for each benchmark function before splitting them between training, validation, and testing. For the benchmark functions that can be defined in dimension higher than two, Ackley and Keane, we impose a dimension of four.
@@ B.2 Additionnal figures
Compared to the normalized results, McCormick is extremely polarized: some models have low nominal test errors while others have bigger ones. While the WaDiRo-SCNN performs well in Experiments A and B, it sees a medium performance in Experiment C. This is hidden by the extremely poor performance of the deep FNN in Experiment C. Figure 2 shows that McCormick has linear tendencies in the center of its domain but steep increases on its edges that linear models fail to capture.
@@
One interesting takeaway from Figure 7 is that WaDiRo-SCNN, at its worst performance, is on par with linear regression models. We do see the standard SCNN formulations have a poorer performance than the linear models on Keane and the deep FNN has consistently worse performances in the corrupted setting from Experiment C.
@@ B.3 Computational time
The tests presented in Figure 8 were conducted on a computer with 64GB of RAM DDR5 and an AMD Ryzen 9 7950x3d CPU. The training was performed with the open-source solver Clarabel and was formulated in Python with cvxpy. As we can see, more neurons lead to smaller training errors while a greater dataset size increases the training error. Indeed, sampling more data points increases the difficulty of fitting them all correctly. As for the training time, to keep the computational time low, there is a tradeoff between the maximal number of neurons and the size of the dataset. This is expected partly because we do not yet have a GPU acceleration method for this training procedure. We also remark that a more direct implementation, e.g., by using a lower-level programming language and avoiding an interface like cvxpy for the constraint implementation, could also diminish the training time. As such, comparing these running times with methods implemented in C++ would be unfair.
@@ C.1 Hyperparameters and setting
{tblr}
colspec=m2.5cm—lll, hline1,8 = 1.5pt, hline2,3,4,5,6,7=1pt Model & Hyperparameter Description Possible values
@@
orm used in Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
orm used in Wasserstein metric $[\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
Loss function used $[\mathrm{e}^{0.1},\mathrm{e}^{7}]$
@@
Noise level of the white noise kernel $[\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$[\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$[\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$[\mathrm{e}^{0.001},\mathrm{e}^{2}]$

### v3

@@ 8.2.2 Numerical Results
We observe that the WaDiRo-SCNN performs on par with the GP as both models seem adequate to model nonlinear patterns. Still, its performance exceeds that of the GP as it has the lower MAE between the two on 10 buildings out of 16 while having similar RMSEs. Note that trained GPs may only converge to local optima. We observe that the linear SVR has the worst performance of all models as it has a consistently lower performance than the WaDiRo linear regression. We observe no significant performance gap between the unconstrained and physics-constrained WaDiRo-SCNNs. We also remark, as shown in Table 3, that the WaDiRo-PCSCNN has the lowest number of constraint violations throughout the experiment and that the WaDiRo-SCNN has the second lowest amount with yet 59.12% more. Additional figures of the absolute errors and the test prediction plots, are presented in Appendix C for completeness.
@@ 8.3.2 Numerical results
Experiment results are presented in Figures 5 and 6. We observe some general tendencies across the two datasets and the different models. The minimum and maximum of each scale are identified in the top left corner of each subfigure. First, the lowest test errors are achieved for high neuron counts and low regularization. Instability seems inversely correlated with the size of the regularization parameter. Thus, there seems to be a trade-off between precision and stability. We note that some neurons count are anomalously less stable than others, e.g., 6 neurons in Figure 5. We believe that this is related to the population of sampling vectors used because the same outlying patterns emerge similarly for all models.
The experiment on the energy dataset demonstrates that neural networks can be highly unstable with lower regularization. Indeed, with the added effect of the scaling, $L_{\varepsilon}$ reaches the millions. For a large number of neurons and an intermediate regularization parameter, we note a favourable balance between low test errors and model stability.
@@ 9 Closing remarks
Acknowledgements
This work was possible thanks to funding from the Fonds de recherche du Québec (FRQ), the Natural Sciences and Engineering Research Council of Canada (NSERC), Mitacs Accelerate, and Hilo by Hydro-Québec under scholarships CGRS-M and B1X, as well as Grants RGPIN-2023-04235 and IT35303 & IT38517. Special thanks to Salma Naccache and Bertrand Scherrer for their active support and the enriching discussions, as well as Steve Boursiquot, Ahmed Abdellatif, and Odile Noël from Hilo for making this project possible.
@@ A Proof of Proposition 1
$h_{\bm{\beta}}^{*}(\bm{\theta})<+\infty\iff\sup_{\hat{\mathbf{z}}:\hat{\mathbf{z}}^{\top}\bm{\beta}\geq 0}\{\bm{\theta}^{\top}\hat{\mathbf{z}}-{\bm{\beta}}^{\top}\hat{\mathbf{z}}\}<+\infty\ \ \text{and}\ \sup_{\hat{\mathbf{z}}:\hat{\mathbf{z}}^{\top}\bm{\beta}\leq 0}\{\bm{\theta}^{\top}\hat{\mathbf{z}}+{\bm{\beta}}^{\top}\hat{\mathbf{z}}\}$$\infty$.
@@
$\max\ (\bm{\theta}-{\bm{\beta}})^{\top}\hat{\mathbf{z}}$ $\xrightarrow{\text{dual}}\$$\min\ 0\cdot\lambda_{A}$ (Dual-A)
$\text{s.t.}\ \hat{\mathbf{z}}^{\top}\bm{\beta}\geq 0$ $\text{s.t.}\ {\bm{\beta}}\lambda_{A}=\bm{\theta}-{\bm{\beta}},\ \lambda_{A}\leq 0,$
@@
$\max(\bm{\theta}+{\bm{\beta}})^{\top}\hat{\mathbf{z}}$ $\xrightarrow{\text{dual}}\$$\min\ 0\cdot\lambda_{B}$ (Dual-B)
$\text{s.t. }\ \hat{\mathbf{z}}^{\top}\bm{\beta}\leq 0$ $\text{s.t.}\ {\bm{\beta}}\lambda_{B}=\bm{\theta}+{\bm{\beta}},\ \lambda_{B}\geq 0.$
@@
$\exists\lambda_{A}\leq 0,\quad\text{s.t.}\quad$ ${\bm{\beta}}\lambda_{A}=\bm{\theta}-{\bm{\beta}}$
$\exists\lambda_{B}\geq 0,\quad\text{s.t.}\quad$ ${\bm{\beta}}\lambda_{B}=\bm{\theta}+{\bm{\beta}}.$
@@
$\min_{\bm{\nu},\bm{\omega}}\quad$ $\epsilon a\$$+\frac{1}{N}\sum_{j=1}^{N}c_{j}$
@@
$-{{\beta}}_{i}$ $\leq a\quad\forall i\in\llbracket 2Pd+1\rrbracket$
${{\beta}}^{\top}\hat{\mathbf{z}}_{j}$ $\leq c_{j}\quad\forall j\in\llbracket N\rrbracket$
@@
[…]
@@
[…]
@@ B.1 Hyperparameters and setting
Model Hyperparameter Description Possible values
@@
Norm of the Wasserstein metric $\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Regularization framework $\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Norm of the Wasserstein metric $\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
Regularization framework $\mathrm{e}^{-8},\mathrm{e}^{1}]$
@@
dropout_p Size of each batch
@@
We sample 2000 data points for each benchmark function before splitting them between training, validation, and testing. For the benchmark functions that can be defined in dimensions higher than two, Ackley and Keane, we impose a dimension of four.
@@ B.2 Additionnal figures
Compared to the normalized results, McCormick is extremely polarized: some models have low nominal test errors while others have bigger ones. While the WaDiRo-SCNN performs well in Experiments A and B, it sees a medium performance in Experiment C. This is hidden by the extremely poor performance of the deep FNN in Experiment C. Figure 2 shows that McCormick has linear tendencies in the centre of its domain but steep increases on its edges that linear models fail to capture.
@@
One interesting takeaway from Figure 7 is that WaDiRo-SCNN, at its worst performance, is on par with linear regression models. We do see that the standard SCNN formulations have a poorer performance than the linear models on Keane, and the deep FNN has consistently worse performances in the corrupted setting from Experiment C.
@@ B.3 Computational time
The tests presented in Figure 8 were conducted on a computer with 64GB of RAM DDR5 and an AMD Ryzen 9 7950x3d CPU. The training was performed with the open-source solver Clarabel and was formulated in Python with cvxpy. As we can see, more neurons lead to smaller training errors, while a greater dataset size increases the training error. Indeed, sampling more data points increases the difficulty of fitting them all correctly.
To keep the training time low, there is a trade-off between the maximal number of neurons and the size of the dataset. This is expected partly because we do not yet have a GPU acceleration method for this training procedure. We also remark that a more direct implementation, e.g., by using a lower-level programming language such as C++ and avoiding an interface like cvxpy to formulate the optimization problem, could also greatly diminish the training time. Our implementation is thus relatively inefficient. As such, we do not compare the training times with other models from the literature compared in Section 8, because most of them are developed in highly computationally optimized or GPU-enabled libraries, i.e., scikit-learn and pytorch. The comparison would be unfair when dealing with large datasets without the proper computation time optimization of our method. This is left for future work as discussed in Section 9. Nonetheless, as can be seen in Figure 8, our method can be relatively fast under some dimensionality constraints.
@@ C.1 Hyperparameters and setting
Model Hyperparameter Description Possible values
@@
Norm used in Wasserstein metric $\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
Norm used in Wasserstein metric $\mathrm{e}^{-6},\mathrm{e}^{1}]$
@@
Loss function used $\mathrm{e}^{0.1},\mathrm{e}^{7}]$
@@
Noise level of the white noise kernel $\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$\mathrm{e}^{0.1},\mathrm{e}^{5}]$
$\mathrm{e}^{0.001},\mathrm{e}^{2}]$